OpenStudy (anonymous):

Write as a sum of partial fractions. x/(3x-2)(2x+1)

OpenStudy (anonymous):

\[\begin{align*}\frac{x}{(3x-2)(2x+1)}&=\frac{A}{3x-2}+\frac{B}{2x+1}\\ x&=A(2x+1)+B(3x-2)\\ x&=(2A+3B)x+A-2B \end{align*}\] This gives you the system \[\begin{cases}2A+3B=1\\ A-2B=0\end{cases}\] Solve for \(A\) and \(B\).